OpenStudy (anonymous):

Find the quotient. If possible, rename the quotient as a mixed number. 2/5÷2/8=

OpenStudy (anonymous):

8/5

OpenStudy (anonymous):

no is not

OpenStudy (anonymous):

Ans is 8/5. When expressed as mixed number, it becomes,1 whole 3/5.

OpenStudy (anonymous):

yes quotient is 1

OpenStudy (anonymous):

there we go :) ^^

OpenStudy (anonymous):

2/5x8/2=16/10=8/5 16/10=1+6/10 mixed number 8/5=1.6=1+6/10 mixed number